Dad is 42 years old and his son is 10. In how many years will Dad be 3 times as old as his son?
Answer by mananth(16946)

let the number of years after which dad is 3 times be x
42+x = 3(10+x)
42+x = 30+3x
42-30=3x-x
12=2x
x=6 after 6 years.
...
m.ananth@hotmail.ca